What is point-slope form of an equation of the line through the points (–3, 8) and (7, 6). Use (–3, 8) as the point for the equation.
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Use the two-point form of an equation of a line to derive an equation for the desired line:



where and are the coordinates of the given points.

Once you have made the appropriate substitutions and done the indicated arithmetic, rearrange your equation into slope-intercept () form by solving for

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

What is point-slope form of an equation of the line through the points (–3, 8) and (7, 6). Use (–3, 8) as the point for the equation.
----------------------
slope = (6-8)/(7--3) = -2/10 = -1/5
----
an Form:
8 = (-1/5)(-3)+b
---
b = 8 - 3/5
---
b = 37/5
----
Equation:
y = (-1/5)x+(37/5)
=======================
Cheers,
Stan H.